Two spherical balls of mass 10 kg each are placed 10 cm apart. Find the gravitational force of attraction between them.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. Gravitational force of attraction, F = GMm/r^2 = 6.67 * 10^-11 * 10 * 10/(0.2)^2 = 6.67 * 10^-7 N